how do you do3/4z=12 explain how to do it
Answer by stanbon(75887)

how do you do3/4z=12 explain how to do it
---
You want to know the value of "Z".
----
In your problem z is multiplied by (3/4)
---
You need to multiply 3/4 by 4/3 to get z isolated:
---
But the equality sign means you also have to multiply the right
side by 4/3.
---
(4/3)(3/4)z = (4/3)12
---
Then:
z = (4/3)12
Multiplying you get:
z = 16
==================
Cheers,
Stan H.